### v2

# from Answers to Reports

Abstract
As an embodiment of intelligence evolution toward interconnected architectures, Deep Research Agents (DRAs) systematically exhibit the capabilities in task decomposition, cross-source retrieval, multi-stage reasoning, information integration, and structured output, which markedly enhance performance on complex and open-ended tasks. However, existing benchmarks remain deficient in evaluation dimensions, response format, and scoring mechanisms, limiting their effectiveness in assessing such agents. This paper introduces Dr. Bench, a multidimensional evaluation framework tailored to DRAs and long-form report-style responses. The benchmark comprises 214 expert-curated challenging tasks across 10 broad domains, each accompanied by manually constructed reference bundles to support composite evaluation. This framework incorporates metrics for semantic quality, topical focus, and retrieval trustworthiness, enabling a comprehensive evaluation of long reports generated by DRAs. Extensive experimentation confirms the superior performance of mainstream DRAs over web-search-tool-augmented reasoning models, yet reveals considerable scope for further improvement. This study provides a robust foundation for capability assessment, architectural refinement, and paradigm advancement of DRAs.
Yang Yao1, 2Yixu Wang1, 3Yuxuan Zhang4Yi Lu5Tianle Gu1, 6Lingyu Li1, 7
Dingyi Zhao1, 7Keming Wu6Haozhe Wang8Ping Nie9Yan Teng1, 🖂Yingchun Wang1
1 Shanghai Artificial Intelligence Laboratory 2 The University of Hong Kong 3 Fudan University
4 University of British Columbia 5 University of Toronto 6 Tsinghua University 7 Shanghai Jiao Tong University
8 Hong Kong University of Science and Technology 9 Peking University
yaoyangacademia@outlook.com, tengyan@pjlab.org.cn
1 Introduction
The core architecture of AI paradigms is undergoing a transition from closed, static, parameter-driven Large Language Models (LLMs) to active, cognitive, and interconnected agent systems endowed with external perception and integrative mechanisms (Wang et al., 2024; Zhang et al., 2025b). Amid growing demands for heterogeneous information acquisition and increasing task complexity, Deep Research has gradually emerged as a paradigm of agent systems characterized by cognitive planning and integrative capabilities (Xu and Peng, 2025), which enable autonomous task decomposition, cross-source retrieval, multi-step reasoning, and structured expression, thereby substantially enhancing model adaptability and expressive performance in real-world applications (Huang et al., 2025; Zhang et al., 2025a).
Although Deep Research Agents (DRAs) demonstrate strong task execution capabilities, the current benchmarking framework remains significantly outdated in both design philosophy and coverage scope. First, existing benchmarks predominantly target discrete short-text outputs, such as multiple-choice answers or brief phrases (Wei et al., 2024; Zhou et al., 2025; Ho et al., 2020; Yang et al., 2018), which, although conducive to efficient evaluation and automation, cannot be extended to complex report-style generation tasks and fail to reflect the demands of DRAs for logical inference and linguistic organization. Second, most benchmarks continue to assess isolated competencies, focusing primarily on reasoning or web search, without establishing systematic criteria for evaluating integrated performance. In particular, mechanisms for assessing citation authority, source validity, and semantic drift in long-form outputs remain absent. Mainstream evaluation methods rely either on string matching (Cohen et al., 2025; Monteiro et al., 2024), which fails to capture semantic adequacy, or on similarity scoring with LLMs as judgers (Chen et al., 2025a; Pham et al., 2025), which lacks transparent and verifiable standards and is therefore prone to subjectivity and instability.
To address the limitations inherent in existing evaluations, our study proposes a rigorous benchmark, authored by human experts and specifically designed for DRAs, which targets high-difficulty and high-precision tasks with the goal of systematically assessing the overall performance in report-style long-text generation. On this basis, we develop a multidimensional evaluation framework intending to accurately and stably measure the quality and credibility of generated reports. Our research yields three key contributions:
  1. (1)
    We introduce the D(eep)r(esearch). Bench paired with manually constructed reference bundles comprising query-specific and general-report rubrics, trustworthy sources, focus-anchor and focus-deviation keywords. Covering 214 challenging report-oriented entries across 10 broad domains, Dr. Bench rigorously and efficiently enables broad and granular characterization of DRA tasks, which addresses existing deficiencies in evaluation dimensions and response formats.
  2. (2)
    We develop a systematic and multidimensional evaluation framework for long-form report-style outputs, which captures key processes of DRAs such as task reasoning, information retrieval, content synthesis, and structured articulation. By jointly modeling semantic quality, topical focus, and retrieval trustworthiness, our framework overcomes limitations of conventional methods and demonstrates high transferability to long-text generations beyond DRAs.
  3. (3)
    We conduct large-scale experiments involving five mainstream DRAs, one advanced agent model, and seven reasoning models enhanced with web-search tools. Quantitative results indicate that DRAs consistently outperform tool-augmented models in overall task execution proficiency and report generation quality, while revealing persistent limitations in architectural paradigms and behavioral mechanisms that warrant further refinement and continued development.
2 Related Works
2.1 Deep Research Agents
Early LLMs, such as GPT-3 (Brown et al., 2020) and PaLM (Chowdhery et al., 2023), rely on static training corpora and closed parameter spaces, while their knowledge is entirely derived from the training phase and cannot be updated or supplemented during inference. To overcome the epistemic constraints of static language models, researchers have explored mechanisms for integrating LLMs with external tools, giving rise to the paradigm of Tool-Augmented LLMs, exemplified by GPT-4 (Achiam et al., 2023), Gemini 1.5 (Google, 2024), Claude 3 (Anthropic, 2024), Toolformer (Schick et al., 2023), and Qwen 1.5 Agent (Alibaba, 2024). These models leverage external interfaces such as web browsers to enable dynamic information acquisition and cross-modal perception, reflecting a shift from knowledge encapsulation toward tool-mediated cognition.
Beyond generative and inferential capabilities, DRAs integrate cognitive planning and information fusion to support end-to-end workflows for complex and open-ended tasks. Their functionality encompasses task recognition, multi-stage decomposition, heterogeneous retrieval, cross-source aggregation, and structured report generation, emphasizing system-level autonomy and procedural integrity. This paradigm is broadly applicable to academic research, policy analysis, technology evaluation, and market intelligence, while remaining applicable to daily-life scenarios for general users. Representative systems include open-source DRAs like Tongyi DeepResearch (Alibaba, 2025) and commercial platforms such as Grok Deep Search (xAI, 2025), Sonar Deep Research (Perplexity, 2025), and o3 Deep Research (OpenAI, 2025), which implement full-spectrum research pipelines. These systems exemplify a shift from static knowledge encapsulation to cognitively extended intelligence, advancing LLMs toward agentic architectures with strategic reasoning capabilities.
2.2 Existing Benchmarks
With the growing adoption of Tool-Augmented LLMs, both academia and industry have increasingly focused on their performance in web-search tasks. Existing benchmarks, including GAIA (Mialon et al., 2023), WebWalker (Wu et al., 2025), BrowseComp (Wei et al., 2025), WideSearch (Wong et al., 2025), BrowseComp-Plus (Chen et al., 2025b), and Deep Research Bench (Bosse et al., 2025), primarily evaluate LLMs using closed-form queries that require verifiable short answers to facilitate automated scoring and alignment. However, they lack coverage of key behaviors such as task decomposition, cross-source retrieval, and structured synthesis, and provide limited assessment of content hierarchy, discourse structure, and information integration. A rough comparison of existing benchmarks can be found in Appendix F. In addition, most rely on surface-level matching or similarity metrics, such as Exact Match, BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and BERTScore (Zhang et al., 2019), which struggle to capture semantic depth and structural fidelity, thereby limiting their effectiveness in evaluating the true capabilities of DRAs.
Contemporaneous work has begun to explore evaluation methods tailored to report-style outputs. DeepResearch Bench (Du et al., 2025) is the first to assess reference answers alignment and retrieval. However, it depends heavily on static reference reports, making it difficult to accommodate evolving query expectations. Its automated rubrics lack contextual sensitivity and focus on generic surface, failing to reflect human preferences for report quality and structure. Its retrieval evaluation emphasizes consistency between statements and cited links but overlooks the credibility of sources. Benchmarks such as ResearchQA (Yifei et al., 2025), DeepResearch Arena (Wan et al., 2025), and ReportBench (Li et al., 2025) constructed 21K, 10K, and 0.6K academic tasks respectively. These benchmarks also rely on automatically generated rubrics that exhibit limited stability and interpretability, which undermines their reliability for high-precision judgment. Additionally, large-scale benchmarks increase evaluation costs, especially for expensive DRAs, reducing their practical utility.
Overall, existing benchmarks fail to rigorously and comprehensively evaluate report-style long-form outputs in alignment with human expectations. They lack precise rubrics and trustworthy references, making it difficult to systematically characterize the full capabilities of DRAs.
3 Dr. Bench
3.1 Domains
The Dr. Bench was meticulously constructed and repeatedly validated by human experts, comprising 214 high-complexity entries across diverse domains. It is designed to challenge existing DRAs in task understanding, decomposition, execution, and aggregation. Based on semantic relevance, entries are systematically categorized into ten broad domains, with detailed distribution provided in Figure 1 and Appendix A. The dataset exhibits a high degree of professionalism and diversity in its design, effectively simulating complex and dynamic queries encountered in real-world scenarios, and demonstrating broad coverage and strong representativeness.
Figure 1: The proportion of Dr. Bench entries across domains.
3.2 Components
Each entry consists of a query instruction paired with a reference bundle designed to support performance evaluation. The bundle comprises five core modules: Query-Specific Rubrics (QSRs), General-Report Rubrics (GRRs), Trustworthy-Source Links (TSLs), Focus-Anchor Keywords (FAKs), and Focus-Deviation Keywords (FDKs), each of which corresponds to a distinct capability that a DRA is expected to demonstrate.
3.2.1 Query
As the core input to DRAs, queries are designed to elicit structured long-form reports rather than traditional short, discrete answers, while embodying diversity and representativeness by (1) covering topics from ancient history to contemporary events, including both long-term technological evolution and short-term dynamic shifts; (2) encompassing major countries and regions, with systematically designed cross-regional comparative tasks; (3) involving diverse disciplines, with many tasks intentionally designed for interdisciplinary integration; (4) entailing structured texts across various styles including academic writing, data-driven analysis, technical deconstruction, strategic planning, policy evaluation, and event reconstruction; (5) incorporating multi-level causal chains and cross-source information fusion; and (6) including both quantitative and qualitative analysis.
To enhance consistency and reproducibility, Dr. Bench systematically incorporates spatiotemporal robustness into the design of queries and rubrics. For fact-based tasks, each query explicitly defines temporal and geographic boundaries to mitigate external fluctuations in response content. For open-ended tasks, rubrics emphasize logical structure and reasoning validity to ensure generalizability and adaptability. All queries adopt the instruction “write a report/essay/…” to unify output format and reinforce task orientation. Overall, the query design achieves a high degree of semantic complexity, task diversity, and evaluative operability.
3.2.2 Query-Specific Rubrics
QSRs are designed to evaluate the task completion quality. Each QSR is custom-built by experts based on corresponding query, reflecting human expectations and evaluative preferences regarding factual accuracy and logical validity. Scoring follows a clearly defined binary (Yes/No) or ternary (Partial) scheme to ensure consistency and operational clarity. Each query is equipped with at least 8 QSRs, with a total score of 30, assigned in accordance with task importance.
QSRs are deeply embedded within the semantic structure of each task, offering high alignment and diagnostic precision. Their design spans core dimensions such as information coverage, mechanism explanation, structured expression, semantic precision, source verification, evidence organization, heterogeneity analysis, methodological transparency, temporal logic, and interdisciplinary integration. Appendix G outlines the core dimensions of QSRs design together with their corresponding descriptions. The QSRs provide both theoretical grounding and practical guidance for evaluating DRAs performance, while also laying a structural foundation for future automated scoring systems.
Figure 2: Pipeline for benchmark construction and overview of the evaluation framework.
3.2.3 General-Report Rubrics
GRRs assess the quality of structured expression. Independent of specific queries, GRRs evaluate reports from a general perspective using binary judgment across 7 key dimensions, namely structural organization, logical clarity & expression, informational coverage & content depth, citation quality & source credibility, originality & insight, data usage & analytical rigor, and formatting consistency.
GRRs constitute a necessary, objective, and universally applicable set of standards that delineate the fundamental human expectations for report-style responses: a qualified report must not only accomplish the assigned task but also satisfy general quality requirements such as structural completeness, logical coherence, and insightful argumentation. Comprising 48 rubrics with a total score of 73, the design emphasizes generalizability, normative clarity, and expressive strength, thereby providing a unified quality benchmark for performance evaluation across tasks and models, and establishing a comprehensive evaluation framework grounded in the dual pillars of content and expression, as outlined in Appendix B.
3.2.4 Trustworthy-Source Links
TSLs serve as indicators of the trustworthiness of heterogeneous retrieval and cross-source aggregation. They are designated by experts in variable quantities, and consist of durable, stable, and reliable website links that are authoritative, official, accessible, and contain original information necessary to answer the query. Subjective or non-primary sources such as forums and blogs are excluded. Each link is precisely anchored to the specific page containing the target information, ensuring accuracy, verifiability, and confidence in information acquisition.
To prevent TSLs expiration, we (1) defined stability and longevity requirements for queries and TSLs during data construction, mandating that TSLs originate from authoritative, official, and primary sources (e.g., government or academic institutions) that are reliably accessible and resistant to change, thereby ensuring stability from the outset; (2) introduced hostname matching at the metric mechanism to reduce dependence on specific pages, and incorporated hyperparameters in TrustworthyBoost, enabling full-match stability to be adjusted across evaluation targets; and (3) established long-term maintenance through regular link validity checks (e.g., quarterly) and the planned use of snapshots as permanent references.
3.2.5 Focus-Anchor Keywords
FAKs are used to evaluate thematic focus during cross-source content aggregation. Each set of FAKs is specified by expert designers for the given query and consists of 5 semantically stable core terms, while avoiding superficial phrasing from the query itself. FAKs serve to assess the thematic focus and key point coverage of generated content, enabling effective evaluation in terms of semantic consistency and analytical depth.
3.2.6 Focus-Deviation Keywords
FDKs are used to evaluate the degree of thematic drift. Each set consists of five terms that are prone to triggering topic divergence. Their presence typically indicates that the generated content has deviated from the original query focus, resulting in reduced semantic coherence and increased informational noise. In high-cost DRA processes, such deviations lead to unnecessary consumption of resources, ultimately compromising overall performance.
For example, when a query focuses on “men’s football matches”, the term “women” would divert the report off-topic. Similarly, when the query concerns the “electric vehicles”, references to “fuel cars” or “motorcycles” are classified as FDK. During data construction, these terms were annotated by experts and validated through multiple reviews to ensure their consistency and appropriateness.
3.3 Construction Pipeline
To ensure the difficulty, quality, and semantic validity, Dr. Bench adopts a multi-stage review pipeline for systematic verification and refinement of expert-generated entries. This process integrates manual design, machine auditing, and cross-validation to establish a highly granular framework for data generation and quality control, as illustrated in the upper half of Figure 2.
In the initial stage, experts design diverse data units based on the unified Construction Guide and their own expertise within predefined domains. The units then undergo LLMs auditing to detect semantic inconsistencies, logical errors, and factual inaccuracies. During the 1st round of manual review, QSRs are verified for validity, with rubric criteria refined. For binary items with intermediate cases, a clearly defined “Partial” option is introduced to enhance scoring granularity and consistency. After preliminary filtering, queries and QSRs are standardized and rewritten to ensure stylistic and structural correctness. TSLs and Keywords are also supplemented and revised at this stage. The revised entries then undergo LLM-based difficulty testing, followed by the 2nd round of manual review, which focuses on evaluating QSRs from observers’ perspective and assessing the anchoring effectiveness of keywords. Each link is individually verified for accessibility and authority. The 3rd round of cross-review conducts a comprehensive quality check, covering domain classification, formatting standards, and structural integrity. Examples can be found in Appendix C.
This multi-phase mechanism significantly improves accuracy, consistency, and reproducibility, while also reducing subjective bias and annotation errors, providing a robust foundation for model evaluation.
4 Evaluation Framework
Built on semantic quality, topical focus, and retrieval trustworthiness, a structured and multidimensional evaluation framework is proposed for report-style generation tasks, featuring an integrated scoring system inspired by principles from statistics and operations. It prioritizes transparency, interpretability, and practical utility for robust evaluation of complex tasks, as illustrated in the lower half of Figure 2.
4.1 Semantic Quality
Semantic quality evaluates the overall performance of response reports in terms of task completion and general quality. It integrates scores from both QSRs and GRRs. Drawing on the Weighted Average Method (WAM) from Multi-Attribute Decision Making (MADM), the metric applies ratio normalization to both scores and assigns weighting coefficients $\alpha$ and $\beta$, where $\alpha+\beta=1$, to construct a fused model of multidimensional quality signals. The calculation formula of Quality $\in[0,1]$ is as follows:
Quality $=\alpha\cdot\mathcal{N}_{\text{Ratio}}\left[\sum_{i=1}^{N}\text{QSR}_{\text{score}}^{(i)}\right]+\beta\cdot\mathcal{N}_{\text{Ratio}}\left[\sum_{j=1}^{M}\text{GRR}_{\text{score}}^{(j)}\right]$
where $\text{QSR}_{\text{score}}^{(i)}$ denotes the $i$-th QSR score; $\text{GRR}_{\text{score}}^{(j)}$ refers to the $j$-th GRR score; $N$ and $M$ represent the number of QSRs and GRRs respectively; $\mathcal{N}_{\text{Ratio}}[\cdot]=\cdot\,/\,\sum_{i}\mathrm{FullScore}(r_{i})$ denotes the ratio normalization; $\alpha$ and $\beta$ are weighting parameters reflecting the relative importance in the overall quality assessment.
4.2 Topical Focus
Topical focus is evaluated through the SemanticDrift metric, which jointly considers the absence of FAKs and the misuse of FDKs to measure the degree of thematic deviation.
$\text{FAK}_{\text{Drift}}\in[0,1]$ quantifies the omission of core keywords. The frequency of each keyword in the report is scaled by its expected value $\epsilon$, and a minimum function is introduced to implement a threshold control. A FAK is penalized when its frequency falls below $\epsilon$, thereby enforcing the requirement for substantive keyword presence and ensuring that the score reflects semantic completeness rather than accidental occurrence. Drawing on the TF × IDF paradigm from information retrieval, $\text{FAK}_{\text{Drift}}$ is constructed as the product of frequency component and semantic relevance:
$\text{FAK}_{\text{Drift}}=1-\frac{1}{K}\sum_{k=1}^{K}\left[\min\left(\frac{\text{freq}^{(k)}}{\epsilon_{+}},1\right)\times\mathcal{N}_{\text{Ratio}}(\text{rele}^{(k)})\right]$
where $\text{freq}^{(k)}$ is frequency of the $k$-th FAK; $\text{rele}^{(k)}\in\{1,2,3,4,5\}$ is the semantic relevance assessed by LLMs; $K$ represents the number of FAKs; $\epsilon_{+}$ is the expectation scaling factor. A higher $\text{FAK}_{\text{Drift}}$ value indicates weaker thematic focus and insufficient coverage of core concepts in the report. Similarly, $\text{FDK}_{\text{Drift}}\in[0,1]$, which quantifies the intensity of thematic distraction, is defined as follows:
$\text{FDK}_{\text{Drift}}=\frac{1}{L}\sum_{l=1}^{L}\left[\min\left(\frac{\text{freq}^{(l)}}{\epsilon_{-}},1\right)\times\mathcal{N}_{\text{Ratio}}(\text{rele}^{(l)})\right]$
where $\text{freq}^{(l)}$ is frequency of the $l$-th FDK; $\text{rele}^{(l)}\in\{1,2,3,4,5\}$ is the relevance score; $L$ represents the number of FDKs; $\epsilon_{-}$ is the scaling factor. A higher $\text{FDK}_{\text{Drift}}$ indicates stronger thematic deviation. The SemanticDrift is a weighted combination of $\text{FAK}_{\text{Drift}}$ and $\text{FDK}_{\text{Drift}}$:
$\text{SemanticDrift}=\lambda\cdot\text{FAK}_{\text{Drift}}+\mu\cdot\text{FDK}_{\text{Drift}}\in[0,1]$
where $\lambda+\mu=1$, reflecting the relative importance. SemanticDrift reflects the degree of thematic deviation in the generated report, with higher values indicating weaker alignment to the intended topic and lower values suggesting stronger semantic focus and consistency.
4.3 Retrieval Trustworthiness
Retrieval Trustworthiness evaluates the credibility of external information retrieval and usage in response reports. Modeling based on the hit rate of TSLs and using confidence enhancement mechanism inspired by multiplicative fusion in Bayesian updating, this approach transforms match rates into multiplicative scoring factors to increase the evaluative weight of citation quality. Specifically, matches are categorized into full matches and hostname matches, where $\text{Rate}_{\text{full\,hit}}$ serves as a recall-like metric capturing the precise coverage of provided TSLs, and $\text{Rate}_{\text{host\,hit}}$ reflects the proportion of generalized mentions whose annotation links share the same source domains as the recommended references. These two are assigned different weights and combined to compute the TrustworthyBoost $\in[1,1+\eta]$ as:
$\text{TBoost}=1+\eta\cdot\bigg[\ \theta\cdot\underbrace{\left(\frac{\text{match}_{\text{full}}}{S}\right)}_{\text{Rate}_{\text{full\,hit}}}+\kappa\cdot\underbrace{\left(\frac{\text{match}_{\text{host}}}{T+1}\right)}_{\text{Rate}_{\text{host\,hit}}}\ \bigg]$
where $\text{match}_{\text{full}}$ indicates the number that have been exactly matched in TSLs; $\text{match}_{\text{host}}$ refers to the number of annotations sharing the same hostname as the TSLs; $S$ and $T$ denote the sizes of the TSLs and the annotations respectively; $\theta$ and $\kappa$ represent the weights for full and hostname matches respectively, with $\theta+\kappa=1$. The coefficient $\eta$ controls the magnitude of the confidence boost, thereby preventing excessive inflation due to high confidence and avoiding complete nullification when confidence is low. This mechanism preserves scoring stability while enhancing evaluative sensitivity to verifiability and source reliability, thereby improving the responsiveness to external evidence signals.
4.4 Integrated Scoring Framework
The integrated metric $\in[0,120]$ adopts a multiplicative weighting model to enable multidimensional evaluation of report-style generation tasks.
IntegratedScore $=\text{Quality}\times(1-\text{SDrift})\times\text{TBoost}$
Here, Quality assesses the structural integrity and content quality of the report, SemanticDrift reflects the degree of thematic deviation and is transformed into a positive factor via $1-\text{SemanticDrift}$, and TrustworthyBoost enhances the credibility weight based on authoritative link coverage. This design logic penalizes semantic drift and rewards external support, producing a normalized score on a 100-point scale.
Algorithm 1 Multidimensional Evaluation Framework
1:  for entry in entries do
2:  Extract query, QSRs, GRRs, TSLs, FAKs, FDKs
3:  Extract report, annotations, token_usage
4:  # Semantic Quality
5:  $\text{QSR}_{\text{scores}}\leftarrow\sum\text{QSR}_{\text{score}}$ from LLM-Judger over QSRs
6:  $\text{GRR}_{\text{scores}}\leftarrow\sum\text{GRR}_{\text{score}}$ from LLM-Judger over GRRs
7:  $\text{Quality}\leftarrow\alpha\cdot\mathcal{N}_{\text{Ratio}}(\text{QSR}_{\text{scores}})+\beta\cdot\mathcal{N}_{\text{Ratio}}(\text{GRR}_{\text{scores}})$
8:  # Retrieval Trustworthiness
9:  $\text{Rate}_{\text{full\,hit}}\leftarrow$ $\text{Match}_{\text{full}}/|\texttt{TSLs}|$
10:  $\text{Rate}_{\text{host\,hit}}\leftarrow$ $(\text{Match}_{\text{host}}-\text{Match}_{\text{full}})/|\texttt{annotations}+1|$
11:  $\text{TrustworthyBoost}\leftarrow 1+\eta\cdot\left(\theta\cdot\text{Rate}_{\text{full\,hit}}+\kappa\cdot\text{Rate}_{\text{host\,hit}}\right)$
12:  # Topical Focus
13:  for FAK in FAKs do
14:   $\text{relevance}\leftarrow$ LLM-Judger(FAK, report)
15:   $\text{frequency}\leftarrow$ count of FAK in report
16:   $\text{FAK}_{\text{score}}\leftarrow\min(\text{frequency}/\epsilon_{+},1)\cdot\mathcal{N}_{\text{Ratio}}(\text{relevance})$
17:  end for
18:  $\text{FAK}_{\text{Drift}}\leftarrow 1-\sum\text{FAK}_{\text{score}}/|\texttt{FAKs}|$
19:  Perform same procedure for FDKs to compute $\text{FDK}_{\text{Drift}}$
20:  $\text{SemanticDrift}\leftarrow\lambda\cdot\text{FAK}_{\text{Drift}}+\mu\cdot\text{FDK}_{\text{Drift}}$
21:  # Integrated Scoring Framework
22:  $\text{InteScore}\leftarrow\text{Quality}\cdot(1-\text{SDrift})\cdot\text{TBoost}\cdot 100$
23:  $\text{ContriPerToken}\leftarrow\text{InteScore}/({\text{token}_{\text{total}}-\text{token}_{\text{input}}})$
24:  end for
25:  # Aggregate Metrics
26:  $\overline{\text{InteScore}}\leftarrow$ average of InteScore over entries
27:  $\overline{\text{ContriPerToken}}\leftarrow$ average of ContriPerToken over entries
As shown in Algorithm 1, the framework addresses limitations in traditional evaluation methods when applied to DRAs. It offers strong scalability and transferability, making it broadly applicable to performance assessment of structured long-text generation tasks, particularly those involving tool-augmented systems.
Table 1: Results of the main experiment, with the highest score in bold and the second-highest underlined in each column.
Models Quality $\uparrow$ 11-SDrift $\uparrow$ TBoost $\uparrow$ InteScore $\uparrow$ Usage C/Token $\downarrow$
Qwen-deep-research 0.6348 0.5248 1.0288 34.6480 9258 0.0100
Sonar-deep-research 0.6184 0.5271 1.0238 33.4668 8254 0.0043
o3-deep-research-2025-06-26 0.6176 0.5184 1.0171 32.9004 25038 0.0014
Kimi-K2-0905-preview 0.6707 0.4671 1.0153 32.0651 2079 0.0164
Grok-4-0709-search 0.6130 0.4890 1.0283 31.3490 3012 0.0112
Gemini-2.5-pro 0.5506 0.4856 1.0130 27.3364 5446 0.0072
o4-mini-deep-research-2025-06-26 0.5666 0.4803 1.0203 28.0391 18640 0.0016
GPT-5-2025-08-07 0.5560 0.4593 1.0383 27.3312 7006 0.0045
GPT-4o-search-preview-2025-03-11 0.4945 0.4496 1.0073 22.5645 1005 0.0247
GPT-4.1-2025-04-14 0.4762 0.4694 1.0027 22.4382 1252 0.0194
Claude-opus-4-1-20250805 0.4559 0.4674 1.0202 22.0047 2267 0.0101
Claude-sonnet-4-20250514 0.4491 0.4735 1.0184 21.7235 2267 0.0097
Claude-3-7-sonnet-20250219 0.3996 0.4737 1.0148 19.3415 2327 0.0084
4.5 Additional Metrics
To enable a more comprehensive assessment of DRAs ability, we design a set of supplementary metrics based on accessible response metadata to characterize model performance and efficiency in real-world execution, including token consumption, number of reasoning steps, and the volume of links involved during retrieval.
$\text{ContributionPerToken}=\frac{\text{IntegratedScore}}{token_{\text{\ total}}-token_{\text{\ input}}}$
To evaluate cost-effectiveness under limited resources, the Contribution/Token metric evaluates model efficiency based on actual token expenditure in reasoning and generation, effectively measuring the information density per token. When long texts maintain high information density, efficiency scores can be high; they are truly lowered only when the text contains substantial redundancy or off-topic content, which is precisely what we aim to detect. This metric is designed to encourage models to produce efficient outputs.
$\text{RetrievalIndex}=\frac{num_{\text{annotated}}}{num_{\text{retrieved}}+1}\in[0,1]$
RetrievalIndex evaluates the filtering and aggregation capability of DRAs during the retrieval process. It is defined as the ratio between the number of annotations ultimately adopted in the report and the total number of links retrieved during the search phase, reflecting the model’s ability to effectively distill valuable content from large-scale information. A lower index indicates stronger selectivity and aggregation, suggesting that the DRA can extract more relevant and informative content from redundant sources, thereby enhancing the specificity and density of the generated output.
5 Experiments
5.1 Overview
We evaluated a total of thirteen models under Dr. Bench, including five DRAs (o3-deep-research-2025-06-
26
, qwen-deep-research, sonar-deep-resea-
rch
, grok-4-0709-search, and o4-mini-deep-
research-2025-06-26
), one advanced agent (kimi-
k2-0905-preview
), and seven web-search-tool-enhanced reasoning models (gemini-2.5-pro, gpt-
5-2025-08-07
, gpt-4o-search-preview-2025- 03-11, gpt-4.1-2025-04-14, claude-opus-4-
1-20250805
, claude-sonnet-4- 20250514, and claude-3-7-sonnet-20250219).
To eliminate the randomness, all models with adjustable temperature were evaluated at $\text{temperature}=0.0$. For non-DRA models without embedded annotations, reports were merged with annotations during evaluation. gpt-4o-20-
24-11-20
independently scored each rubric to evaluate semantic quality. Topical focus was assessed on pure report text without annotations to avoid interference from annotation titles. Related prompts are provided in Appendix D. Retrieval Trustworthiness was computed from pure annotations, excluding parameters, anchors, and duplicates to ensure fair matching. Manual verification was randomly conducted on approximately 35% of the scores judged by LLMs, yielding a 99.3% agreement with human evaluations.
Regarding parameters, $\alpha=\beta=0.5$ were set to balance task completion and general quality; $\lambda=0.7$ and $\mu=0.3$ were configured to emphasize sensitivity to the omission of FAKs; coefficient was set to $\eta=0.2$ to control score inflation and prevent over-amplification; $\theta=0.7$ and $\kappa=0.3$ were set for the contributions of exact citations and generalized mentions. This configuration ensures scoring stability while enhancing sensitivity to semantic relevance, citation accuracy, and task fidelity.
5.2 Leaderboard
Table 1 and Appendix E presents the leaderboard results, with models ranked in descending order of IntegratedScore. In terms of IntegratedScore, Qwen ranked first, demonstrating strong performance across all dimensions. Sonar followed closely, achieving the highest score in topical focus. Notably, Kimi-K2, a Mixture-of-Experts architecture agent with 1T parameters, attained the highest score in the quality, outperforming all DRAs. For topical focus, Sonar, Qwen, and o3 formed the leading cluster. GPT-5 achieved the highest score in citation reliability, indicating strong external support alignment. In terms of resource consumption, o3 and o4-mini averaged 23K and 18K tokens per report respectively, placing them lowest in contribution efficiency.
Overall, DRAs exhibited stable semantic quality and control in structured report generation, while web-search-tool-augmented models showed promise in external support and efficiency. These results validate the Dr. Bench’s ability to differentiate model capabilities across multiple dimensions, providing a robust empirical foundation for future optimization of DRAs.
Figure 3: The proportion of quality share across models.
Figures 3 and 10 summarize the role of Quality scores and cross-domain performance in the overall IntegratedScore. While Kimi-K2 achieves a strong Quality score, its relative weaknesses in credibility and attention reduce the impact of this advantage, whereas models such as Qwen, Sonar, and o3 exhibit a more balanced interplay between Quality and multiplicative factors, yielding stronger integrated performance. Across domains, all top models perform notably well in domain 03 (Sports & Competitions) and domain 10 (Health & Medicine), while maintaining relatively balanced results elsewhere. In Quality, Kimi-K2 leads, followed by Qwen and Sonar; however, in SemanticDrift and TrustworthyBoost, Kimi-K2 shows clear disadvantages, whereas the others remain comparatively consistent. Detailed analysis is provided in Appendix E.
5.3 Supplementary Dimensions
Additionally, four OpenAI models yielded richer response metadata, revealing strategic differences in task execution and tool usage beyond the three core dimensions, and offering valuable supplement into multidimensional evaluation.
Table 2: Average inference and retrieval times on OpenAI models.
Models ReasonTimes SearchTimes
GPT-4.1 Unavailable 0.3925
GPT-5 12.9346 10.6729
o3-dr 55.4333 16.1000
o4-mini-dr 63.9860 26.5093
As shown in Table 2, GPT-4.1 showed minimal retrieval activity, with an average of only 0.39 times. In contrast, both o4-mini-dr and o3-dr displayed intensive inference and retrieval patterns, suggesting more complex reasoning chains and information acquisition strategies. In terms of retrieval efficiency, o3-dr slightly exceeded in total retrieved links and annotations, while o4-mini-dr achieved slightly lower RetrievalIndex, reflecting stronger filtering and citation precision, as shown in Table 3.
Table 3: RetrievalIndex of o3-dr and o4-mini-dr.
Models $\text{{num}}_{\text{{retr}}}$ $\text{{num}}_{\text{{anno}}}$ RIndex $\downarrow$
o4-mini-dr 14.4583 7.7986 0.5520
o3-dr 15.2744 8.7561 0.5804
6 Discussions
Two systemic limitations emerged during evaluation that underscore key challenges in the design of DRAs. First, instability in invocation behavior was observed in models such as o3 and o4-mini, which exhibited substantial variance in reasoning time across repeated queries. This suggests a lack of internal constraints governing search frequency and direction, resulting in non-convergent retrieval paths and inconsistent response behavior. Second, semantic decomposition occasionally produced sub-queries in non-English languages with incoherent semantics, despite all tasks being in English. These outputs were misaligned with task intent and unintelligible to human evaluators, thereby impairing retrieval precision and relevance.
These limitations reflect two fundamental trade-offs in DRAs development. The efficiency—quality trade-off highlights the tension between high-quality reasoning and computational cost, with current models often incurring excessive token usage and latency. Addressing this requires adaptive control over search depth and token allocation. Meanwhile, the decomposition—coherence trade-off reveals that while modular query breakdown enhances coverage, it risks semantic fragmentation and intent drift. Future architectures must reconcile decomposition benefits with coherent multi-stage reasoning to ensure consistent task fidelity.
7 Conclusions
In this paper, we present the Dr. Bench and the multidimensional evaluation framework that systematically assesses the performance and capabilities of DRAs. By leveraging challenging queries across diverse thematic domains and high-quality reference bundles, our framework enables rigorous evaluation of report-style outputs along axes of semantic quality, topical focus, and retrieval trustworthiness. Empirical results show that contemporary DRAs substantially outperform conventional tool-augmented models in complex task scenarios, while also exposing key limitations and trade-offs. These insights elucidate current challenges in DRA design and lay the groundwork for advancing DRAs as efficient, stable, and interpretable intelligent agents.
Appendix A Taxonomy of Domains
Figure 4: Taxonomy of domains.
Figure 4 and Table 4 present the domain taxonomy underlying the benchmark corpus. The classification scheme comprises ten principal domains, delineated according to thematic relevance: Academia & Research, News & Current Affairs, Sports & Competitions, Commonsense & Education, Law & Politics, Business & Finance, Technology Intelligence, Environment & Sustainability, History & Social Sciences, and Health & Medicine. Entries that do not align with the predefined domains are assigned to the residual class Unclassified, in order to preserve high-quality data while maintaining diversity.
Table 4: Taxonomy of domains with simplified descriptions.
Code Domain Description Count
00 Unclassified Cannot be categorized 3
01 Academia & Research Academic trends, research methods, etc. 23
02 News & Current Affairs International news, regional hotspots, etc. 18
03 Sports & Competitions Olympics, World Cup, athlete data, match reports, etc. 15
04 Commonsense & Education Common facts, educational resources, etc. 15
05 Law & Politics Legal texts, policy updates, international relations, etc. 24
06 Business & Finance Market analysis, investment strategies, etc. 35
07 Technology Intelligence Artificial intelligence, tech trends, etc. 17
08 Environment & Sustainability Climate change, environmental policies, ecosystems, etc. 12
09 History & Social Sciences History, philosophy, sociology, psychology, etc. 33
10 Health & Medicine Disease research, public health, nutrition knowledge, etc. 19
Summary 214
It is evident that Business & Finance and History & Social Sciences account for a relatively larger proportion of the benchmark. Nevertheless, the overall distribution remains broadly balanced, exhibiting both thematic richness and representational breadth.
Appendix B General-Report Rubrics
Figure 5: Detailed criteria for General-Report Rubrics.
Appendix C Examples of Entries
To illustrate the precision and rigor of our benchmark, we present four representative entries: ID 04216, 07001, 08004, and 10236. These entries span distinct domains and query types and serve as concrete examples of structural completeness, rubric coverage, and citation fidelity.
Figure 6: Example ID 07001 of benchmark entries.
Figure 7: Example ID 04216 of benchmark entries.
Figure 8: Example ID 08004 of benchmark entries.
Figure 9: Example ID 10236 of benchmark entries.
Appendix D Scoring Prompts
As part of the rubric evaluation process, the following prompt was provided to guide scoring consistency and rule adherence.
You are a scoring evaluator tasked with assessing the quality of a report generated by a deep research model. You will be provided with: 1. A report text
2. An evaluation rule containing specific scoring criteria and allowed score values Your task is to: - Carefully read the report
- Evaluate it strictly against the given rule
- Assign a score based only on the score values defined in the rule Scoring instructions: - Only use the score values explicitly listed in the rule
- Do not invent intermediate scores or alternative formats
- Your output must begin with the score in square brackets [], followed by a one-sentence reason Output format example: [0] No citations were provided, which violates the requirement. [2] The report fully meets the requirement with clear and relevant details. Be objective and consistent. Focus on clarity, completeness, relevance, and adherence to the rule.
Report text: {report}
Rule: {rubric}
As part of the keyword relevance evaluation stage, the following prompt was provided to guide consistent scoring.
You are a scoring evaluator tasked with assessing the relevance of a specific keyword within a research report. You will be provided with: 1. A report text
2. A keyword to evaluate Your task is to: - Read the report carefully
- Judge how semantically relevant the keyword is to the report
- Consider not just frequency, but depth of discussion, thematic importance, and contextual integration Use the following 5-point relevance scale: (5) Extremely Relevant: The keyword is a central theme of the report; It appears multiple times and is discussed in depth; The report’s main arguments or findings revolve around it; (4) Highly Relevant: The keyword is a major topic; It appears more than once and is clearly explained or referenced; contributes directly to the report’s purpose; (3) Moderately Relevant: The keyword is mentioned but not emphasized; It may appear once or twice; It supports the report contextually but is not a focus; (2) Slightly Relevant: The keyword is briefly mentioned; It has little impact on the report’s core content; It may be incidental or peripheral; (1) Not Relevant: The keyword does not appear in the report; Or it appears in a way that is unrelated to the report’s topic. Output format example: [4] The keyword "QUIC" is referenced multiple times in the report, particularly in the context of protocol evolution and RFC publication. While not the sole focus, it is clearly a major topic. Be objective and consistent. Focus on clarity, completeness, relevance, and adherence to the rule.
Report text: {report}
Keyword: {keyword}
Appendix E Supplementary Experimental Observations
E.1 Quality Share Proportion
Figure 3 illustrates the contribution of Quality scores to the overall IntegratedScore. As a core metric among the three evaluation dimensions, Quality’s share proportion provides a direct indication of each DRA’s capacity for textual precision, thematic focus, and control over external referencing. Notably, although Kimi-K2 achieves a prominent score in the Quality dimension, its relative deficiencies in credibility and attention metrics diminish the extent to which its quality advantage is reflected in the overall score. In contrast, models such as Qwen, Sonar, and o3 demonstrate a more balanced allocation between Quality and multiplicative factors, exhibiting great integrative performance while leaving room for further optimization.
E.2 Evaluation Across Domains
Table 5 presents a fine-grained evaluation of various models across distinct domains. In this table, D denotes the domain, and M refers to the evaluation metrics, including QUA (Quality), SDR (11-SemanticDrift), TBO (TrustworthyBoost), and ITS (IntegratedScore). Each column corresponds to a model represented by an abbreviation: QWE (qwen-deep-research), SON (sonar-deep-research), O3D (o3-deep-research-2025-06-26), KIM (kimi-k2-0905-preview), GRO (grok-4-0709-search), GEM (gemini-2.5-pro), O4D (o4-mini-deep-research-2025-06-26), GT5 (gpt-5-2025-08-07), G4O (gpt-4o-search-preview-2025-03-11), G41 (gpt-4.1-2025-04-14), OPU (claude-opus-4-1-20250805), SO4 (claude-sonnet-4-20250514), and S37 (claude-3-7-sonnet-20250219). This table is designed to highlight model-specific performance variations across multiple dimensions, providing a structured basis for subsequent analysis.
The distribution of ITS values reveals substantial variation in model performance across domains. In particular, QWE, SON, and O3D consistently achieve higher scores across multiple domains. Their ITS values are notably elevated in domain 03, where they reach 41.27, 38.10, and 40.23 respectively, and in domain 10, with corresponding scores of 39.07, 38.86, and 37.68. These results indicate their stable advantages in these specific contexts. In contrast, models such as S37, SO4, and OPU tend to exhibit lower scores across most domains, reflecting a performance floor. This distribution suggests that models vary in their domain sensitivity and adaptability, with certain systems demonstrating enhanced integrative capabilities under domain-specific conditions.
Table 5: Comparative evaluation across domains.
D M QWE SON O3D KIM GRO GEM O4D GT5 G4O G41 OPU SO4 S37
01 QUA 0.6273 0.6090 0.6340 0.6705 0.6465 0.5774 0.5690 0.5439 0.5013 0.4690 0.4288 0.4362 0.3800
SDR 0.5184 0.5340 0.5457 0.4664 0.5089 0.4873 0.4826 0.4491 0.4667 0.4715 0.4584 0.4616 0.4765
TBO 1.0209 1.0357 1.0304 1.0373 1.0423 1.0161 1.0351 1.0476 1.0102 1.0021 1.0337 1.0245 1.0209
ITS 33.3293 33.8947 36.0683 32.7436 34.4484 29.1322 28.9792 26.8605 23.6181 22.3786 21.3256 20.8999 18.9583
02 QUA 0.6190 0.6331 0.6108 0.6432 0.6077 0.5730 0.5572 0.4939 0.5160 0.4937 0.4457 0.4562 0.3887
SDR 0.5112 0.5282 0.4943 0.4324 0.4760 0.4841 0.4668 0.4276 0.4668 0.4722 0.4508 0.4673 0.4751
TBO 1.0146 1.0131 1.0210 1.0070 1.0222 1.0099 1.0158 1.0126 1.0063 1.0055 1.0175 1.0149 1.0096
ITS 32.0267 34.0071 31.8369 28.4567 30.1611 28.1130 26.3594 22.3089 24.4736 23.4507 20.9942 21.8205 18.6347
03 QUA 0.6725 0.6198 0.6462 0.7139 0.6414 0.5513 0.5438 0.5349 0.5211 0.4551 0.4928 0.4579 0.4307
SDR 0.5909 0.6008 0.6053 0.5756 0.5651 0.5545 0.5743 0.4812 0.5316 0.5283 0.5265 0.5719 0.5875
TBO 1.0441 1.0248 1.0195 1.0158 1.0160 1.0062 1.0220 1.0307 1.0065 1.0000 1.0149 1.0151 1.0147
ITS 41.2741 38.1028 40.2341 41.8823 37.3244 31.4358 31.6925 27.7716 28.2281 24.3922 26.6635 26.5932 25.6712
04 QUA 0.6123 0.6006 0.6465 0.6368 0.5766 0.5329 0.5488 0.5608 0.4589 0.4530 0.4394 0.4250 0.3690
SDR 0.5155 0.4971 0.5094 0.4789 0.4628 0.4833 0.4807 0.4560 0.3981 0.4401 0.4460 0.4287 0.4259
TBO 1.0108 1.0218 1.0142 1.0132 1.0322 1.0186 1.0163 1.0213 1.0126 1.0065 1.0137 1.0108 1.0219
ITS 31.6576 30.1116 33.7925 31.1660 27.8452 26.0326 26.9029 27.1555 19.0310 20.2277 19.8213 18.3432 16.1287
05 QUA 0.5891 0.6012 0.5678 0.6463 0.6066 0.5390 0.5694 0.5517 0.4804 0.4664 0.4435 0.4295 0.3900
SDR 0.5031 0.5142 0.4673 0.4249 0.4728 0.4748 0.4649 0.4461 0.4268 0.4623 0.4676 0.4548 0.4534
TBO 1.0153 1.0270 1.0061 1.0075 1.0288 1.0108 1.0262 1.0301 1.0052 1.0054 1.0187 1.0106 1.0085
ITS 30.8795 32.4488 27.2325 28.1336 30.3342 26.6297 27.8097 25.7082 20.3981 21.7361 21.0149 19.8689 17.8701
06 QUA 0.6257 0.6143 0.6039 0.6840 0.6051 0.5214 0.5319 0.5450 0.4816 0.4567 0.4591 0.4520 0.4036
SDR 0.5320 0.5099 0.5205 0.4856 0.4994 0.4701 0.4617 0.4649 0.4382 0.4671 0.4613 0.4739 0.4685
TBO 1.0490 1.0201 1.0106 1.0094 1.0347 1.0079 1.0112 1.0235 1.0035 1.0035 1.0139 1.0137 1.0103
ITS 35.9545 32.1937 32.0019 33.8587 31.9281 24.9650 25.6285 26.8070 21.4294 21.5077 21.8287 21.7307 19.1475
07 QUA 0.6221 0.6042 0.6104 0.6358 0.5898 0.5416 0.6181 0.5799 0.5026 0.4566 0.4275 0.4408 0.3859
SDR 0.5118 0.5251 0.5044 0.4735 0.4845 0.4972 0.4958 0.4735 0.4569 0.4729 0.4702 0.4699 0.4624
TBO 1.0577 1.0290 1.0215 1.0262 1.0170 1.0093 1.0266 1.0529 1.0127 1.0014 1.0235 1.0165 1.0131
ITS 33.4141 32.6757 31.5103 30.8186 29.5214 27.2947 31.2982 29.2195 23.2293 21.7192 21.0743 21.2224 18.4858
08 QUA 0.6741 0.6338 0.5994 0.6610 0.5945 0.5004 0.5095 0.5806 0.4744 0.4685 0.4504 0.4532 0.3981
SDR 0.5093 0.4833 0.4762 0.4198 0.4145 0.4422 0.4350 0.4252 0.4007 0.4380 0.4085 0.4292 0.4048
TBO 1.0179 1.0175 1.0136 1.0060 1.0219 1.0107 1.0124 1.0158 1.0053 1.0022 1.0153 1.0221 1.0119
ITS 35.1035 30.7007 28.8171 27.6535 25.0956 22.2533 23.2147 25.4842 18.9471 20.2026 18.4119 19.7139 16.2288
09 QUA 0.6460 0.6245 0.6149 0.6841 0.6019 0.5558 0.5814 0.5700 0.4832 0.5037 0.4657 0.4650 0.4182
SDR 0.5108 0.5152 0.5182 0.4428 0.4752 0.4771 0.4678 0.4622 0.4336 0.4454 0.4616 0.4704 0.4672
TBO 1.0082 1.0176 1.0091 1.0143 1.0225 1.0128 1.0187 1.0501 1.0053 1.0000 1.0270 1.0230 1.0181
ITS 33.5468 32.3533 31.8783 30.6664 29.6514 26.7788 27.3007 28.2353 21.0317 22.0717 21.8632 22.1229 19.6543
10 QUA 0.6666 0.6447 0.6639 0.7129 0.6504 0.6012 0.6218 0.6074 0.5407 0.5155 0.4942 0.4598 0.4148
SDR 0.5514 0.5765 0.5453 0.4967 0.5093 0.5021 0.4982 0.4973 0.4913 0.5101 0.5156 0.5091 0.5149
TBO 1.0507 1.0344 1.0379 1.0148 1.0372 1.0293 1.0190 1.0923 1.0103 1.0019 1.0200 1.0321 1.0185
ITS 39.0697 38.8577 37.6762 36.0924 34.9180 31.0504 31.7808 33.4642 27.0223 26.4282 26.0997 24.2348 22.0003
MIX QUA 0.6348 0.6184 0.6176 0.6707 0.6130 0.5506 0.5666 0.5560 0.4945 0.4762 0.4559 0.4491 0.3996
SDR 0.5248 0.5271 0.5184 0.4671 0.4890 0.4856 0.4803 0.4593 0.4496 0.4694 0.4674 0.4735 0.4737
TBO 1.0288 1.0238 1.0171 1.0153 1.0283 1.0130 1.0203 1.0383 1.0073 1.0027 1.0202 1.0184 1.0148
ITS 34.6480 33.4668 32.9004 32.0651 31.3490 27.3364 28.0391 27.3312 22.5645 22.4382 22.0047 21.7235 19.3415
Figure 10: Radar charts of top models’ performance across domains based on multiple metrics.
Figure 10 presents radar charts illustrating the cross-domain performance of the top five models, with axes rescaled for clarity. In the ITS panel, all models exhibit notably strong performance in domain 03 and domain 10, while maintaining relatively balanced results across other domains. This suggests that the models align most closely with human expectations in the areas of Sports & Competitions as well as Health & Medicine. In the QUA panel, KIM demonstrates a clear advantage, followed by QWE and SON. Conversely, in the SDR and TBO panels, KIM shows a marked disadvantage, whereas the remaining models display comparatively consistent performance.
Appendix F Comparison with Existing Benchmarks
Table 6 presents a comparison of our benchmark with existing ones across key dimensions. Task Type specifies whether the task is open-ended or closed and whether the target response takes the form of a short string or a long report. Evaluation Dimensions indicate which metrics are applied and which aspects of model behavior are examined. Construction method describes whether the benchmark is primarily built through LLM-based generation or through human annotation. Scale refers to the size of the dataset in terms of the number of instances. Target System specifies whether the benchmark is designed for LLMs equipped with web-search tools or for DRAs. Finally, Evaluation Criteria clarify whether performance is judged by exact matching, LLM-based criteria, or expert rubric assessment.
Table 6: Comparison of Benchmarks for Tool-Augmented LLMs
Benchmarks Task Type EvalDim Constru Scale System EvalCriteria
WebWalker  (Wu et al., 2025) Closed/String Accuracy LLMs 680 LLMs Match
BrowseComp-Plus (Chen et al., 2025b) Closed/String Accuracy LLMs 830 DRAs Match
GAIA (Mialon et al., 2023) Closed/String Accuracy Human 466 LLMs Match
BrowseComp (Wei et al., 2025) Closed/String Accuracy Human 1266 LLMs Match
WideSearch (Wong et al., 2025) Closed/String F1 Score Human 200 LLMs Match
Deep Research Bench (Bosse et al., 2025) Closed/String Recall, F1 Score Human 89 LLMs Match
ResearchQA (Du et al., 2025, Concurrent Work) Open/Report Quality LLMs 21K DRAs LLMs Criteria
ReportBench (Li et al., 2025, Concurrent Work) Open/Report Quality LLMs 678 DRAs LLMs Criteria
DeepResearch Arena (Wan et al., 2025, Concurrent Work) Open/Report Quality LLMs 10K DRAs LLMs Criteria
DeepResearch Bench (Du et al., 2025, Concurrent Work) Open/Report Quality Human 100 DRAs LLMs Criteria
Dr. Bench (ours) Open/Report Quality, Semantics, Retrieval Human 214 DRAs Experts Rubrics, Keywords, Links
Dr. Bench is the only benchmark that combines quality, semantic adequacy, and retrieval credibility in evaluating report-style outputs. Unlike large-scale auto-generated datasets, it is human-authored with expert rubrics, anchors, and trustworthy links, ensuring transparency and reproducibility. Its 214 carefully curated tasks balance coverage with precision, making it uniquely suited for systematic and fine-grained assessment of Deep Research Agents. Our evaluation dimensions and methodology differ substantially from prior work, with more qualitative and fine-grained details provided in Section 2.2.
Appendix G Core Dimensions of QSRs
Within the evaluation framework for structured long-text tasks, a critical criterion for assessing the effectiveness and representativeness of QSRs is whether their design demonstrates discriminative capacity in judging task completeness and aligns with human user expectations. QSRs constitute a dimension-based framework tailored to specific queries, intended to measure the completeness of reports. The design rationale of QSRs considers, though is not limited to, the core dimensions outlined in Table 7.
Table 7: Core dimensions with descriptions of QSRs
Dimensions Descriptions
Information Coverage Whether the report fully covers key facts, events, clauses, persons, institutions, or geographic units required by the query, ensuring no essential information is missing.
Technical / Mechanism Explanation Whether the report accurately explains mechanisms, protocols, causal chains, policy tools, or theoretical frameworks involved in the query, reflecting mastery of core logic.
Structural Expression Whether the report generates structured forms such as tables, timelines, maps, matrices, or lists to support organization and reproducibility of the task.
Semantic Precision Whether the report defines key terms, distinguishes concepts, annotates units and time, and avoids ambiguity, ensuring clarity and consistency of expression.
Source Verification Whether the report cites authoritative texts, identifiers, database versions, retrieval times, links, or snapshots to support traceability and credibility of claims.
Evidence Organization Whether the report employs multi-source cross-validation, independent source pairing, counterfactual baselines, placebo tests, or robustness checks to enhance reliability.
Heterogeneity and Comparative Analysis Whether the report identifies and analyzes differences across regions, groups, time periods, or institutions, supporting comparative and explanatory power.
Methodological Transparency Whether the report clearly explains analytical methods, identification strategies, variable definitions, and data processing steps, ensuring consistency and reproducibility.
Social Impact and Distribution Effects Whether the report analyzes distributional impacts of policies or events on different groups, communities, industries, or ecosystems, reflecting social dimensions.
Historical Evolution and Temporal Logic Whether the report establishes event chains, policy evolution paths, time windows, or lag mechanisms, supporting historical and dynamic analysis.
Interdisciplinary Integration Whether the report integrates perspectives from multiple disciplines (e.g., law, economics, technology, society, philosophy) to support complexity and explanatory depth.